OpenStudy (anonymous):

maximum value of x3y3 + 3 x*y when x+y = 8

OpenStudy (cwrw238):

i think this will be maximum when xy is a maximum To find x when xy is maimum: let z = xy x + y = 8 y = 8-x z = x(8 - x) z' = 8 - 2x x = 4 xy = 4*4 = 16 maximum value is 4^3*4^3 + 3*4*4 = 4144

OpenStudy (cwrw238):

* z' = 8 - 2x for max/min 8 - 2x = 0 giving x = 4 maximum value as second derivative is -2 (negative0

OpenStudy (anonymous):

Another method, using Lagrange multipliers. We want to maximize \(x^3y^3+3xy\) subject to \(x+y=8\). Basically we let \(f(x,y)=x^3y^3+3xy\) and \(g(x,y)=x+y=8\), and we have \[\nabla f(x,y)=\lambda \nabla g(x,y)\] which gives the system \[\begin{cases} 3x^2y^3+3y=\lambda(1+0)=\lambda\\ 3x^3y^2+3x=\lambda(0+1)=\lambda\\ x+y=8 \end{cases}\] Setting the first two equations equal to each other: \[\begin{align*}3x^2y^3+3y&=3x^3y^2+3x\\ x^2y^3+y&=x^3y^2+x\\ y\left(x^2y^2+1\right)&=x\left(x^2y^2+1\right)\\ y&=x\end{align*}\] Substituting into the second equation, \(x+y=2x=8\) gives \(x=4\), and hence \(y=4\).